OpenStudy (anonymous):

sove using the substitution method x+y=-2 y-2x=1

OpenStudy (anonymous):

simplify the 2nd equation into y = 2x+1 then plug it into equation one so it becomes x+(2x+1)=-2 the combine like terms and bring variables to one side. 3x + 1 = -2, 3x = -3, x = -1. then plug the x into any of the 2 equations to get y. -1 + y = -2, y = -1